OpenStudy (anonymous):

PLEASSE HELP!! (DNT UNDERSTAND)!!! What is an estimate of the solution of the equation 6n + 3 = 2? us a table

OpenStudy (anonymous):

We know that n=-1/6

OpenStudy (anonymous):

6n + 3 = 2 6n = 2-3 6n = -1 n = -1/6

OpenStudy (anonymous):

I don't know how you would use a table for this problem since there is only one variable n.